OpenStudy (anonymous):

dentify the correct slope and y intercept of the equation 3x - 3y = 4.

OpenStudy (stamp):

3x - 3y = 4 Solve for y = mx + b 3 (x - y) = 4 x - y = 4/3 x - 4/3 = y y = mx + b m = 1, b = -4/3